OpenStudy (anonymous):

Use the foil method to find the product (2x^5-7)(x^9-7)=

OpenStudy (campbell_st):

 First \[2x^5 \times x^9 = 2x^{14}\] Outside \[2x^5 \times - 7 = -14x^5\] Inside \[-7 \times x^9 = -7x^9\] Last \[-7 \times -7 = 49\] combine the terms \[(2x^2 - 7)(x^9 - 7) = 2x^{14} - 14x^5 - 7x^9 + 49\]